# literature.bib

### v1

@@ 1 Introduction
Discrete adjoints are an established methodology to obtain derivatives of simulation codes, e. g. computational fluid dynamic (CFD) solvers, with numerous applications such as design optimization [VerstraeteMM2017, HeMMM2018], uncertainty quantification [HuK2019], parameter identification [AurouxG2017, LaussOSN2018], or adaptive mesh refinement [KastF2013]. This involves differentiating the discretized problem’s solution procedure, where automatic differentiation (AD) has proven its value as a tool that provides accurate, consistent, and maintainable derivatives with bounded performance costs [GriewankW2008, Naumann2011]. Examples for applications of AD to CFD codes include [BischofEtAl1992, ShermanEtAl1996, MuellerC2005, BischofEtAl2006, TowaraN2013, TowaraSN2015].
@@
With computational resources becoming more powerful and easily available, it is nowadays common to apply simulations to large-scale problems in the order of 100 million degrees of freedom distributed across large numbers of compute nodes. To retain scaling efficiency, simulation codes typically need to be adapted to characteristics of the HPC architecture. With increasing numbers of CPU cores per node, it becomes worthwhile to transition to a hierarchical parallelization approach [Rabenseifner2009] and complement distributed memory parallelism (often implemented via MPI [MPI2023]) with shared memory parallelism, for example via OpenMP [OpenMP52]. For discrete adjoints, this poses the challenge to adapt the AD approaches accordingly. For example, [HueckelheimHSM2019] discusses a strategy to reintroduce OpenMP parallelism to a serial discrete adjoint code obtained with the source transformation AD tool Tapenade [HascoetP2013] from an OpenMP parallel primal code, with applications to a flow solver. Tapenade was subsequently extended to handle OpenMP parallel loops directly [HueckelheimH2022].
@@
SU2111https://su2code.github.io/ and https://github.com/su2code is an open-source code for multiphysics simulations [EconomonPCTA2016], with applications, e. g., in aeroelasticity [GomesP2022] or conjugate heat transfer [BurghardtEtAl2021]. SU2 features discrete adjoints [AlbringSG2015, AlbringSG2016], using derivatives provided by the operator overloading AD tool CoDiPack [^2] https://www.scicomp.uni-kl.de/software/codi/ and https://github.com/scicompkl/codipack [SagebaumAG2019]. The MeDiPack add-on [^3] https://www.scicomp.uni-kl.de/software/medi/ and https://github.com/scicompkl/medipack contributes capabilities to differentiate MPI-parallel codes and enables MPI-parallel discrete adjoints [SagebaumEtAl2023]. After explorative performance studies and prototype developments towards HPC improvements in the context of SU2 and initial improvements of the SU2 code [EconomonEtAl2015, EconomonEtAl2016], a recent HPC modernization improves the scalability of SU2’s primal solvers in multiple ways [GomesEP2021]. For example, changes to the storage layout of solution variables improve the memory locality, and SIMD types ensure that SU2 benefits from the vectorization hardware of today’s processors. Furthermore, a new layer of OpenMP parallelism complements the existing MPI parallelism in SU2. It enables hybrid parallel execution as displayed in Fig. 1.1. Within a single node (or CPU, or NUMA domain), distributed memory parallelism can be replaced by shared memory parallelism, with perspectives for performance improvements as discussed in the literature [Rabenseifner2009]. For example, fewer MPI processes lead to a reduced communication overhead. At the same time, there is less data duplication across processes, and fewer communication buffers are required, which can reduce per-node memory usage.
@@
In this work, we extend the hybrid parallelism in SU2 to the discrete adjoint solvers, thereby making these advantages available to them as well. As a key step towards enabling hybrid parallel discrete adjoints, we apply OpDiLib [^4] https://www.scicomp.uni-kl.de/software/opdi/ and https://github.com/scicompkl/opdilib, a novel add-on for operator overloading AD tools that enables the differentiation of OpenMP parallel codes [BluehdornSG2023]. The resulting hybrid differentiation approach is visualized in Fig. 1.1. As such, this work also demonstrates the applicability of OpDiLib to a large code base and its performance on large-scale test cases. While our focus is on single-zone problems and SU2’s Reynolds-averaged Navier-Stokes (RANS) solver [EconomonPCTA2016], our insights and improvements are not specific to this setup, so that our work can also serve as a reference both for extending hybrid parallel discrete adjoints to other solvers within SU2 and for applications of OpDiLib to other software packages and AD workflows.
@@
Figure 1.1: OpenMP-MPI hybrid parallel execution of SU2. The hybrid parallelism is reflected in the AD tools that provide the derivatives. CoDiPack and MeDiPack are already applied to SU2. In the course of this work, we apply OpDiLib for the differentiation of OpenMP.
@@
The outline of this paper is as follows. Section 2 provides a brief introduction to AD and an overview over the components of SU2’s AD workflow, as all of these need to be regarded for the transition to hybrid parallel discrete adjoints that we address in detail in Section 3. We describe optimizations of the hybrid parallel discrete adjoint performance in Section 4. In Section 5.1, we assess the single-socket performance of hybrid parallel discrete adjoints in detail and showcase the effectiveness of the optimizations. We study the performance impact of loop parallelization strategies with respect to hybrid parallel discrete adjoints in Section 5.2. Section 5.3 focuses on the performance at large scale. Throughout, we identify promising targets for further performance improvements. We conclude our work in Section 6.
2 Automatic Differentiation Workflow in SU2
@@
Discrete adjoints in SU2 are based on a fixed-point formulation of the primal equations and an interpretation of the primal solution process as a fixed-point iteration [AlbringSG2015, AlbringSG2016]. We summarize the core ideas. Consider the implicit relation
$U=G(U,X)$ (2.1)
@@
with state $U$, parameters of interest $X$ (e. g., mesh coordinates), and an operator $G$ that represents one step of the primal solver. The primal solver in SU2 iterates (2.1) as $U_{n+1}=G(U_{n},X)$ until convergence to the primal solution $U^{\ast}$.
@@
Given an objective function $J$ depending on $U$ and $X$, a Lagrange multiplier approach for the minimization of $J$ with respect to $X$ subject to the constraint (2.1) yields discrete adjoint equations
$\bar{U}$ $=\frac{\partial}{\partial U}J^{\mathrm{T}}(U^{\ast},X)+\frac{\partial}{\partial U}G^{\mathrm{T}}(U^{\ast},X)\bar{U},$ (2.2)
$\bar{X}$ $=\frac{\partial}{\partial X}J^{\mathrm{T}}(U^{\ast},X)+\frac{\partial}{\partial X}G^{\mathrm{T}}(U^{\ast},X)\bar{U},$ (2.3)
@@
where the adjoint solution $\bar{U}$ is a suitably chosen Lagrange multiplier and $\bar{X}$ stands for the gradient of the objective function with respect to the parameters $X$ [AlbringSG2015, AlbringSG2016]. Alternatively, Eqs. 2.2 and 2.3 are obtained by formulating a reverse accumulation procedure for the differentiation of the fixed point (2.1) according to [Christianson1994]. (2.2) corresponds to a fixed point iteration for the computation of $\bar{U}$. It is iterated until convergence to $\bar{U}^{\ast}$, which is then used to compute $\bar{X}$ according to (2.3). Under suitable assumptions, the adjoint fixed-point iteration inherits the convergence properties of the primal fixed-point iteration [Christianson1994].
@@
The reverse mode of AD is capable of computing the transposed-matrix vector products in Eqs. 2.2 and 2.3 without assembling any of the involved matrices explicitly. With this capability, the linear system of equations (LSE)
$\left(\frac{\partial}{\partial U}G^{\mathrm{T}}(U^{\ast},X)-I\right)\bar{U}=-\frac{\partial}{\partial U}J^{\mathrm{T}}(U^{\ast},X)$ (2.4)
@@
for $\bar{U}$ implied by (2.2) can also be solved by other matrix-free approaches such as Krylov methods [GomesP2022_2]. Besides the fixed-point approach, in SU2 it is possible to use GMRES to solve (2.4) [GomesP2022].
@@
We briefly introduce the reverse mode of AD. See, e. g., [GriewankW2008, Naumann2011] for comprehensive introductions to AD. We view a computer program $F$ after evaluation of the control flow as a computational graph that translates inputs $x$ via intermediate variables $v$ to outputs $y$. Nodes correspond to variables and edges represent the data flow. Each node is computed from its predecessors by means of an elementary mathematical operation $\phi$ with known partial derivatives as
$w=\phi(u_{1},u_{2},\ldots)$ (2.5)
@@
where $w$ is an intermediate variable or an output and the $u_{i}$ are intermediate variables or inputs. For each variable $v$, we introduce a corresponding adjoint variable $\bar{v}=\frac{\mathrm{d}y}{\mathrm{d}v}^{\mathrm{T}}\bar{y}$ that represents the components of the derivative of $y$ viewed as a function of $v$ weighted linearly according to user-provided seeds $\bar{y}$. Adjoint variables can be computed in an incremental fashion by reversing the data flow in the computational graph. Starting at the adjoint variables $\bar{y}$ associated with the output variables, each statement (2.5) is complemented by adjoint updates
$\bar{u}_{i}\kern 3.50006pt\text{\small+}\kern-3.50006pt=\frac{\partial}{\partial u_{i}}\phi(u_{1},u_{2},\ldots)\bar{w}$ (2.6)
for all $i$ followed by $\bar{w}=0$. On the level of the whole program, this procedure corresponds to computing
$\bar{x}=\frac{\mathrm{d}}{\mathrm{d}x}F(x)^{\mathrm{T}}\bar{y},$ (2.7)
that is, a transposed-Jacobian vector product such as needed in Eqs. 2.2 and 2.3.
@@
Listing 2.1: Schematic computation type for operator overloading AD.
The reverse mode of AD as described above is provided to SU2 by the operator overloading AD tool CoDiPack [AlbringSG2015, SagebaumAG2019]. As the operator overloading approach relies on exchanging the floating point computation type, SU2 provides a global typedef su2double and uses it consistently as the floating point computation type throughout the code. Non-AD builds define it as double, whereas AD builds set it to a suitable CoDiPack type. The latter can be thought of as a structured type as depicted in LABEL:listing:su2double for which, amongst others, the common mathematical operators such as $+$ or $\cdot$, math library functions such as $\sin$ or $\exp$, and assignment operators such as $=$ are overloaded. The overloads perform the usual primal computations and store their results in the value members of su2double variables. In addition, the overloads record the computational graph in a stack-like structure commonly referred to as tape. As a part of this, they manage (assign) the identifier members of su2double variables [SagebaumBG2020]. The computational graph is recorded in terms of these identifiers. To perform the incremental updates (2.6), the reverse evaluation of the computational graph uses the identifiers to address into CoDiPack’s internal vector of adjoint variables (that contains the $\bar{u}_{i}$ and $\bar{w}$ in (2.6), not to be confused with $\bar{U}$ and $\bar{X}$ in the discrete adjoint equations (2.2) and (2.3)).
We describe the AD workflow in SU2 with a focus on the discrete adjoint approach given by (2.2) and (2.3). First, the primal equation (2.1) is iterated until convergence to $U^{\ast}$ with a non-AD build of SU2 and subsequently loaded in an AD build of SU2. (2.2) assumes a clean separation of all computations into $G$ and $J$, which can be difficult to achieve in practice. For convenience, a primary recording therefore captures the computational graph of the joint evaluation of $G$ and $J$
$U^{\ast}\mapsto\begin{pmatrix}G(U^{\ast},X)\\ J(G(U^{\ast},X),X)\end{pmatrix},$ (2.8)
@@
$\bar{U}_{i+1}=\frac{\partial}{\partial U}\left(J(G(U^{\ast},X),X)\right)^{\mathrm{T}}\cdot 1.0+\frac{\partial}{\partial U}G^{\mathrm{T}}(U^{\ast},X)\bar{U}_{i},$ (2.9)
which is of the form (2.2) with the modified objective $\tilde{J}(U,X)=J(G(U,X),X)$ (note the additional application of $G$). The updated values $\bar{U}_{i+1}$ are found in the adjoint values associated with the input $U^{\ast}$ to $G$. SU2 stores the iterate $\bar{U}_{i}$. For each iteration of (2.9), SU2 copies the current iterate $\bar{U}_{i}$ into CoDiPack’s internal vector of adjoint variables as seeds, evaluates the tape, and extracts the updated values $\bar{U}_{i+1}$. CoDiPack’s vector of adjoint variables is reset to zero in between iterations. After the fixed-point iteration (2.9) has converged to $\bar{U}^{\ast}$, a secondary recording
@@
is performed in the same spirit as (2.8), with the parameters $X$ as inputs, and a single evaluation with respective seeds $\bar{U}^{\ast}$ and 1.0 corresponds to the realization of (2.3) with the same modified objective $\tilde{J}$ as before. In between these two main recordings, SU2 performs an additional evaluation of $J(G(U^{\ast},X),X)$ without recording. As a side effect, this passive evaluation resets the identifier members of all intermediate and output variables to zero, thus resetting the association with adjoint variables. Note that the identifiers of input variables are reset separately, as there are no statements in the course of $J(G(U^{\ast},X),X)$ that overwrite them. Note also that $\bar{X}$ admits the interpretation as adjoint variables
@@
Recording tapes incurs substantial memory costs. Besides improving the accuracy of the derivative computation, a key benefit of the iterative approach described above is hence that it is sufficient to record the computational graph of a single iteration of $G$, as opposed to recording the whole primal iterative process in a black-box fashion. SU2 makes use of further advanced AD techniques to improve the derivative’s accuracy and to reduce the memory consumed by recordings [AlbringSG2016]. $G$ commonly involves the approximate solution of LSEs as an iterative subroutine. The analytic derivative of solving an LSE is known [Giles2008] and involves solving an LSE with the transposed matrix, which can be done explicitly using the same iterative LSE solvers [AlbringSG2016]. This reduces both the differentiation overhead and improves the derivative’s accuracy. In terms of AD, LSE solution procedures are treated as externally differentiated functions, or external functions in short [Kowarz2008]. Finally, SU2 makes heavy use of local preaccumulation, that is, contracting computational subgraphs [GriewankW2008]. If a subgraph’s number of input and output variables is small compared to its number of intermediate variables, it is more memory-efficient to compute and store the full Jacobian of this subgraph instead of storing the Jacobians for the individual nodes. The Jacobians are assembled by tape recordings and evaluations for subgraphs throughout the recording of the overall tape for $G$. Noting that smaller tapes require less time to evaluate and given that a single iteration of $G$ is recorded once and evaluated multiple times in the course of (2.2), the runtime effort of preaccumulating the Jacobians is compensated for by multiple tape evaluations at reduced runtime cost.
3 Enabling Hybrid Parallel Discrete Adjoints in SU2
Our goal of enabling hybrid parallel discrete adjoints in SU2 affects all parts of the AD workflow discussed in Section 2. As a key step, we extend the AD capabilities to handle the OpenMP parallel parts in $G$ and $J$, including the advanced AD concepts applied in SU2. We identify and remedy code patterns that do not agree well with hybrid parallel AD, and discuss the involved tradeoffs. We also parallelize the AD workflow itself, in particular the management parts. We properly integrate the new features into the existing code structures and seek to preserve key properties of SU2 such as extensibility and maintainability. While we describe changes made to SU2, many of the observations are general. They hold likewise for enabling hybrid parallel discrete adjoints of other solvers within SU2, and can be transferred to applications of OpDiLib beyond SU2.

### v2

# Hybrid parallel discrete adjoints in SU2

@@ Abstract
keywords:
Automatic differentiation , High-performance computing , Discrete adjoint , OpenMP , OpDiLib , SU2
[…]
[scicomp] organization=Chair for Scientific Computing, University of Kaiserslautern-Landau (RPTU), addressline=Paul-Ehrlich-Straße 34/36, city=Kaiserslautern, postcode=67663, country=Germany
[…]
[su2foundation] organization=SU2 Foundation, addressline=Kluyverweg 1, city=Delft, postcode=2629 HS, country=The Netherlands
@@ 1 Introduction
Discrete adjoints are an established methodology to obtain derivatives of simulation codes, e. g. computational fluid dynamic (CFD) solvers, with numerous applications such as design optimization [1, 2], uncertainty quantification [3], parameter identification [4, 5], or adaptive mesh refinement [6]. This involves differentiating the discretized problem’s solution procedure, where automatic differentiation (AD) has proven its value as a tool that provides accurate, consistent, and maintainable derivatives with bounded performance costs [7, 8]. Examples for applications of AD to CFD codes include [9, 10, 11, 12, 13, 14].
@@
With computational resources becoming more powerful and easily available, it is nowadays common to apply simulations to large-scale problems in the order of 100 million degrees of freedom distributed across large numbers of compute nodes. To retain scaling efficiency, simulation codes typically need to be adapted to characteristics of the HPC architecture. With increasing numbers of CPU cores per node, it becomes worthwhile to transition to a hierarchical parallelization approach [15] and complement distributed memory parallelism (often implemented via MPI [16]) with shared memory parallelism, for example via OpenMP [17]. For discrete adjoints, this poses the challenge to adapt the AD approaches accordingly. For example, [18] discusses a strategy to reintroduce OpenMP parallelism to a serial discrete adjoint code obtained with the source transformation AD tool Tapenade [19] from an OpenMP parallel primal code, with applications to a flow solver. Tapenade was subsequently extended to handle OpenMP parallel loops directly [20].
@@
SU2111https://su2code.github.io/ and https://github.com/su2code (both visited on July 16th, 2024) is an open-source code for multiphysics simulations [21], with applications, e. g., in aeroelasticity [22] or conjugate heat transfer [23]. SU2 features discrete adjoints [24, 25], using derivatives provided by the operator overloading AD tool CoDiPack [^2] https://scicomp.rptu.de/software/codi/ and https://github.com/scicompkl/codipack (both visited on July 16th, 2024) [26]. The MeDiPack add-on [^3] https://scicomp.rptu.de/software/medi/ and https://github.com/scicompkl/medipack (both visited on July 16th, 2024) contributes capabilities to differentiate MPI-parallel codes and enables MPI-parallel discrete adjoints [27]. After explorative performance studies and prototype developments towards HPC improvements in the context of SU2 and initial improvements of the SU2 code [28, 29], a recent HPC modernization improves the scalability of SU2’s primal solvers in multiple ways [30]. For example, changes to the storage layout of solution variables improve the memory locality, and SIMD types ensure that SU2 benefits from the vectorization hardware of today’s processors. Furthermore, a new layer of OpenMP parallelism complements the existing MPI parallelism in SU2. It enables hybrid parallel execution as displayed in Fig. 1. Within a single node (or CPU, or NUMA domain), distributed memory parallelism can be replaced by shared memory parallelism, with perspectives for performance improvements as discussed in the literature [15]. For example, fewer MPI processes lead to a reduced communication overhead. At the same time, there is less data duplication across processes, and fewer communication buffers are required, which can reduce per-node memory usage.
@@
In this work, we extend the hybrid parallelism in SU2 to the discrete adjoint solvers, thereby making these advantages available to them as well. As a key step towards enabling hybrid parallel discrete adjoints, we apply OpDiLib [^4] https://scicomp.rptu.de/software/opdi/ and https://github.com/scicompkl/opdilib (both visited on July 16th, 2024), a novel add-on for operator overloading AD tools that enables the differentiation of OpenMP parallel codes [31]. The resulting hybrid differentiation approach is visualized in Fig. 1. As such, this work also demonstrates the applicability of OpDiLib to a large code base and its performance on large-scale test cases. While our focus is on single-zone problems and SU2’s Reynolds-averaged Navier-Stokes (RANS) solver [21], our insights and improvements are not specific to this setup, so that our work can also serve as a reference both for extending hybrid parallel discrete adjoints to other solvers within SU2 and for applications of OpDiLib to other software packages and AD workflows.
@@
Figure 1: OpenMP-MPI hybrid parallel execution of SU2. The hybrid parallelism is reflected in the AD tools that provide the derivatives. CoDiPack and MeDiPack are already applied to SU2. In the course of this work, we apply OpDiLib for the differentiation of OpenMP.
@@
The outline of this paper is as follows. Section 2 summarizes the background of our work. We provide a brief introduction to discrete adjoints and AD in Section 2.1. Section 2.2 highlights the components of SU2’s AD workflow, as all of these need to be regarded for the transition to hybrid parallel discrete adjoints that we address in detail in Section 3. Section 3.1 focuses on the transition from MPI-parallel to hybrid parallel discrete adjoints. We describe optimizations of the hybrid parallel discrete adjoint performance in Section 3.2. In Section 4, we evaluate the performance characteristics of our implementation by means of various performance studies. In Section 4.1, we assess the single-socket performance of hybrid parallel discrete adjoints in detail and showcase the effectiveness of the optimizations. We study the performance impact of loop parallelization strategies with respect to hybrid parallel discrete adjoints in Section 4.2. Section 4.3 focuses on the performance at large scale. Throughout, we discuss the results and identify promising targets for further performance improvements. We conclude our work in Section 5.
2 Background
2.1 Discrete adjoints and automatic differentiation
@@
Discrete adjoints in SU2 are based on a fixed-point formulation of the primal equations and an interpretation of the primal solution process as a fixed-point iteration [24, 25]. We summarize the core ideas. Consider the implicit relation
$U=G(U,X)$ (1)
@@
with state $U$, parameters of interest $X$ (e. g., mesh coordinates), and an operator $G$ that represents one step of the primal solver. The primal solver in SU2 iterates (1) as $U_{n+1}=G(U_{n},X)$ until convergence to the primal solution $U^{\ast}$.
@@
Given an objective function $J$ depending on $U$ and $X$, a Lagrange multiplier approach for the minimization of $J$ with respect to $X$ subject to the constraint (1) yields discrete adjoint equations
$\bar{U}$ $=\frac{\partial}{\partial U}J^{\mathrm{T}}(U^{\ast},X)+\frac{\partial}{\partial U}G^{\mathrm{T}}(U^{\ast},X)\bar{U},$ (2)
$\bar{X}$ $=\frac{\partial}{\partial X}J^{\mathrm{T}}(U^{\ast},X)+\frac{\partial}{\partial X}G^{\mathrm{T}}(U^{\ast},X)\bar{U},$ (3)
@@
where the adjoint solution $\bar{U}$ is a suitably chosen Lagrange multiplier and $\bar{X}$ stands for the gradient of the objective function with respect to the parameters $X$ [24, 25]. Alternatively, Eqs. 2 and 3 are obtained by formulating a reverse accumulation procedure for the differentiation of the fixed point (1) according to [32]. (2) corresponds to a fixed point iteration for the computation of $\bar{U}$. It is iterated until convergence to $\bar{U}^{\ast}$, which is then used to compute $\bar{X}$ according to (3). Under suitable assumptions, the adjoint fixed-point iteration inherits the convergence properties of the primal fixed-point iteration [32].
@@
The reverse mode of AD is capable of computing the transposed-matrix vector products in Eqs. 2 and 3 without assembling any of the involved matrices explicitly. With this capability, the linear system of equations (LSE)
$\left(\frac{\partial}{\partial U}G^{\mathrm{T}}(U^{\ast},X)-I\right)\bar{U}=-\frac{\partial}{\partial U}J^{\mathrm{T}}(U^{\ast},X)$ (4)
@@
for $\bar{U}$ implied by (2) can also be solved by other matrix-free approaches such as Krylov methods [33]. Besides the fixed-point approach, in SU2 it is possible to use GMRES to solve (4) [22].
@@
We briefly introduce the reverse mode of AD. See, e. g., [7, 8] for comprehensive introductions to AD. We view a computer program $F$ after evaluation of the control flow as a computational graph that translates inputs $x$ via intermediate variables $v$ to outputs $y$. Nodes correspond to variables and edges represent the data flow. Each node is computed from its predecessors by means of an elementary mathematical operation $\phi$ with known partial derivatives as
$w=\phi(u_{1},u_{2},\ldots)$ (5)
@@
where $w$ is an intermediate variable or an output and the $u_{i}$ are intermediate variables or inputs. For each variable $v$, we introduce a corresponding adjoint variable $\bar{v}=\frac{\mathrm{d}y}{\mathrm{d}v}^{\mathrm{T}}\bar{y}$ that represents the components of the derivative of $y$ viewed as a function of $v$ weighted linearly according to user-provided seeds $\bar{y}$. Adjoint variables can be computed in an incremental fashion by reversing the data flow in the computational graph. Starting at the adjoint variables $\bar{y}$ associated with the output variables, each statement (5) is complemented by adjoint updates
$\bar{u}_{i}\kern 3.50006pt\text{\small+}\kern-3.50006pt=\frac{\partial}{\partial u_{i}}\phi(u_{1},u_{2},\ldots)\bar{w}$ (6)
for all $i$. A proper reset of $\bar{w}$ that regards both adjoint variable reuse and the special case of the same adjoint variable appearing on both sides of (6) is achieved as follows: create a temporary copy of $\bar{w}$, reset $\bar{w}=0$, and use the temporary copy to perform the updates (6) [7]. On the level of the whole program, this procedure corresponds to computing
$\bar{x}=\frac{\mathrm{d}}{\mathrm{d}x}F(x)^{\mathrm{T}}\bar{y},$ (7)
that is, a transposed-Jacobian vector product such as needed in Eqs. 2 and 3.
@@
Listing 1: Schematic computation type for operator overloading AD.
The reverse mode of AD as described above is provided to SU2 by the operator overloading AD tool CoDiPack [24, 26]. As the operator overloading approach relies on exchanging the floating point computation type, SU2 provides a global typedef su2double and uses it consistently as the floating point computation type throughout the code. Non-AD builds define it as double, whereas AD builds set it to a suitable CoDiPack type. The latter can be thought of as a structured type as depicted in LABEL:listing:su2double for which, amongst others, the common mathematical operators such as $+$ or $\cdot$, math library functions such as $\sin$ or $\exp$, and assignment operators such as $=$ are overloaded. The overloads perform the usual primal computations and store their results in the value members of su2double variables. In addition, the overloads record the computational graph in a stack-like structure commonly referred to as tape. As a part of this, they manage (assign) the identifier members of su2double variables [34]. The computational graph is recorded in terms of these identifiers. To perform the incremental updates (6), the reverse evaluation of the computational graph uses the identifiers to address into CoDiPack’s internal vector of adjoint variables (that contains the $\bar{u}_{i}$ and $\bar{w}$ in (6), not to be confused with $\bar{U}$ and $\bar{X}$ in the discrete adjoint equations (2) and (3)).
2.2 Automatic differentiation workflow in SU2
We describe the AD workflow in SU2 with a focus on the discrete adjoint approach given by (2) and (3). First, the primal equation (1) is iterated until convergence to $U^{\ast}$ with a non-AD build of SU2 and subsequently loaded in an AD build of SU2. (2) assumes a clean separation of all computations into $G$ and $J$, which can be difficult to achieve in practice. For convenience, a primary recording therefore captures the computational graph of the joint evaluation of $G$ and $J$
$U^{\ast}\mapsto\begin{pmatrix}G(U^{\ast},X)\\ J(G(U^{\ast},X),X)\end{pmatrix},$ (8)
@@ 2.2 Automatic differentiation workflow in SU2
$\bar{U}_{i+1}=\frac{\partial}{\partial U}\left(J(G(U^{\ast},X),X)\right)^{\mathrm{T}}\cdot 1.0+\frac{\partial}{\partial U}G^{\mathrm{T}}(U^{\ast},X)\bar{U}_{i},$ (9)
which is of the form (2) with the modified objective $\tilde{J}(U,X)=J(G(U,X),X)$ (note the additional application of $G$). The updated values $\bar{U}_{i+1}$ are found in the adjoint values associated with the input $U^{\ast}$ to $G$. SU2 stores the iterate $\bar{U}_{i}$. For each iteration of (9), SU2 copies the current iterate $\bar{U}_{i}$ into CoDiPack’s internal vector of adjoint variables as seeds, evaluates the tape, and extracts the updated values $\bar{U}_{i+1}$. CoDiPack’s vector of adjoint variables is reset to zero in between iterations. After the fixed-point iteration (9) has converged to $\bar{U}^{\ast}$, a secondary recording
@@
is performed in the same spirit as (8), with the parameters $X$ as inputs, and a single evaluation with respective seeds $\bar{U}^{\ast}$ and 1.0 corresponds to the realization of (3) with the same modified objective $\tilde{J}$ as before. In between these two main recordings, SU2 performs an additional evaluation of $J(G(U^{\ast},X),X)$ without recording. As a side effect, this passive evaluation resets the identifier members of all intermediate and output variables to zero, thus resetting the association with adjoint variables. Note that the identifiers of input variables are reset separately, as there are no statements in the course of $J(G(U^{\ast},X),X)$ that overwrite them. Note also that $\bar{X}$ admits the interpretation as adjoint variables
@@
Recording tapes incurs substantial memory costs. Besides improving the accuracy of the derivative computation, a key benefit of the iterative approach described above is hence that it is sufficient to record the computational graph of a single iteration of $G$, as opposed to recording the whole primal iterative process in a black-box fashion. SU2 makes use of further advanced AD techniques to improve the derivative’s accuracy and to reduce the memory consumed by recordings [25]. $G$ commonly involves the approximate solution of LSEs as an iterative subroutine. The analytic derivative of solving an LSE is known [35] and involves solving an LSE with the transposed matrix, which can be done explicitly using the same iterative LSE solvers [25]. In terms of AD, LSE solution procedures are treated as externally differentiated functions, or external functions in short [36]. Finally, SU2 makes heavy use of local preaccumulation, that is, contracting computational subgraphs [7]. If a subgraph’s number of input and output variables is small compared to its number of intermediate variables, it is more memory-efficient to compute and store the full Jacobian of this subgraph instead of storing the Jacobians for the individual nodes. The Jacobians are assembled by tape recordings and evaluations for subgraphs throughout the recording of the overall tape for $G$. Noting that smaller tapes require less time to evaluate and given that a single iteration of $G$ is recorded once and evaluated multiple times in the course of (2), the runtime effort of preaccumulating the Jacobians is compensated for by multiple tape evaluations at reduced runtime cost.
3 Methods and implementation
3.1 Enabling hybrid parallel discrete adjoints in SU2
Our goal of enabling hybrid parallel discrete adjoints in SU2 affects all parts of the AD workflow discussed in Section 2.2. As a key step, we extend the AD capabilities to handle the OpenMP parallel parts in $G$ and $J$, including the advanced AD concepts applied in SU2. We identify and remedy code patterns that do not agree well with hybrid parallel AD, and discuss the involved tradeoffs. We also parallelize the AD workflow itself, in particular the management parts. We properly integrate the new features into the existing code structures and seek to preserve key properties of SU2 such as extensibility and maintainability. While we describe changes made to SU2, many of the observations are general. They hold likewise for enabling hybrid parallel discrete adjoints of other solvers within SU2, and can be transferred to applications of OpDiLib beyond SU2.